# In this paper, we show that the constant-dimensional Weisfeiler–Leman algorithm for groups (Brachter & Schweitzer, LICS 2020) can be fruitfully used to improve parallel complexity upper bounds on isomorphism testing for several families of groups. In particular, we show:

### v6

@@ 3.2 Coprime extensions with an $O(1)$-generated complement
Let $K$ be a group of order $n$ such that $K\not\cong G$. We will first show that if there is no action such that $K$ decomposes as $H\ltimes N$, then Spoiler can win with $O(1)$ pebbles and $O(1)$ rounds. Let $f:G\to K$ be the bijection that Duplicator selects. As $N\leq G$, as a subset, is uniquely determined by its orders—it is precisely the set of all elements in $G$ whose orders divide $|N|$—we may assume that $K$ has a normal Hall subgroup of size $|N|$. For first, if for some $n\in N$, $|n|\neq|f(n)|$, Spoiler can pebble $n\mapsto f(n)$ and win immediately. By reversing the roles of $K$ and $G$, it follows that $K$ must have precisely $|N|$ elements whose orders divide $|N|$. Second, suppose that those $|N|$ elements do not form a subgroup. Then there are two elements $x,y\in f(N)$ such that $xy\notin f(N)$. At the first round, Spoiler pebbles $a:=f^{-1}(x)\mapsto x$. Let $f^{\prime}:G\to K$ be the bijection Duplicator selects at the next round. As $K$ has precisely $|N|$ elements of order dividing $|N|$, we may assume that $f^{\prime}(N)=f(N)$ (setwise). Let $b\in N$ s.t. $f^{\prime}(b)=y$. Spoiler pebbles $b\mapsto y$. Now as $N$ is a group, $ab\in N$. However, as $f(a)f^{\prime}(b)\not\in f(N)$, $|ab|\neq|f(a)f^{\prime}(b)|$. So the map $(a,b)\mapsto(x,y)$ does not extend to an isomorphism. Spoiler now wins.
Now we have that $f(N)$ is a subgroup of $K$, and because it is the set of all elements of these orders, it is characteristic and thus normal. Suppose that $f(N)\not\cong N$. We have two cases: either $f(N)$ is not Abelian, or $f(N)$ is Abelian but $N\not\cong f(N)$. Suppose first that $f(N)$ is not Abelian. Let $x\in f(N)$ such that $x\not\in Z(f(N))$, and let $g:=f^{-1}(x)\in N$. Spoiler pebbles $g\mapsto x$. Let $f^{\prime}:G\to K$ be the bijection that Duplicator selects at the next round. We may again assume that $f^{\prime}(N)=f(N)$ (setwise), or Spoiler wins with two additional pebbles and two additional rounds. Now let $y\in f(N)$ such that $[x,y]\neq 1$. Let $h\in G$ such that $f^{\prime}(h)=y$. Spoiler pebbles $h\mapsto y$. Now the map $(g,h)\mapsto(x,y)$ does not extend to an isomorphism, so Spoiler wins. Suppose instead that $f(N)$ is Abelian. As Abelian groups are determined by their orders, we have by the discussion in the first paragraph that Spoiler wins with $2$ pebbles and $2$ rounds.
@@ 4 A “rank” lemma
We say that $C\subseteq G$ is distinguished by the version $J$ pebble game with $k$ pebbles and $r$ rounds if for all $g\in C$ and $g^{\prime}\notin C$, Spoiler can win this pebble game, played on $G$ and $G$, starting from the configuration $g\mapsto g^{\prime}$. In terms of WL colorings, this is equivalent to: after $r$ rounds of $(k+1)$-WL Version $J$ coloring, the colors of all tuples containing elements of $C$ are distinct from the colors of all tuples not containing elements of $C$.
@@
Our primary uses of this lemma in this paper are to show that if $C$ is distinguished by $(k,r)$-WL, then $\langle C\rangle$ is distinguished by $(k+O(1),r+O(\log n))$-WL. However, the preservation of $C$-rank itself, rather than merely the subgroup generated by $C$, seems potentially useful for future applications. In particular, Lem. 4.3 shows that WL can identify verbal subgroups in $O(\log n)$ rounds, provided WL can readily identify each word.
@@
Let $g\in C,g^{\prime}\not\in C$. By assumption, Spoiler can win with $k$ pebbles and $r$ rounds, starting from the configuration $g\mapsto g^{\prime}$. It follows that $\chi_{k,r}((g,\ldots,g))\neq\chi_{k,r}((g^{\prime},\ldots,g^{\prime}))$, where $\chi_{k,r}$ is the coloring computed by $(k,r)$-WL Version $J$. Let $C^{\prime}\subseteq H$ be the set of elements in $H$ whose $\chi_{k,r}$ color is the same as one of the colors in $C$, that is,
$C^{\prime}:=\{h\in H:\chi_{k,r}((h,\ldots,h))\in\{\chi_{k,r}((g,\ldots,g)):g\in C\}\}.$
Then for any bijection $f:G\to H$ chosen by Duplicator, we have that $f(C)=C^{\prime}$, or Spoiler can win with $k$ pebbles in $r$ rounds. We will thus use $\operatorname{rk}(x)$ to denote $\operatorname{rk}_{C}(x)$ if $x\in G$, and $\operatorname{rk}_{C^{\prime}}(x)$ if $x\in H$. We proceed by induction on the rank.
By assumption, rank-1 elements must be sent to rank-1 elements, since $C=\{g\in G:\operatorname{rk}(g)=1\}$, or Spoiler can win with $k$ pebbles in $r$ rounds.
Let $q\geq 1$, and suppose that for all $x\in G$, if $\operatorname{rk}(x)\leq q$, then $\operatorname{rk}(f(x))=\operatorname{rk}(x)$. Suppose $x\in G$ is such that $\operatorname{rk}(x)=q+1$ but $\operatorname{rk}(f(x))\neq\operatorname{rk}(x)$. Since $f$ is a bijection on elements of smaller rank, the only possibility is $\operatorname{rk}(f(x))>q+1$. Spoiler begins by pebbling $x\mapsto f(x)$.
@@
Let $f^{\prime}:G\to H$ be the bijection that Duplicator selects at the next round. Write $x=x_{1}\cdots x_{q+1}$, where for each $i$, $x_{i}\in C$. For $11\leq i\leq j\leq q+$ 1, write $x[i,\ldots,j]:=x_{i}\cdots x_{j}$. We consider the following cases.
@@
  • Case 1: Suppose first that $\operatorname{rk}(y)=\operatorname{rk}(f^{\prime}(y))$ for all $y\in G$ with $\operatorname{rk}(y)\leq q$. In this case, Spoiler pebbles $x[2,\ldots,q+1]\mapsto f^{\prime}(x[2,\ldots,q+1])$. Let $f^{\prime\prime}:G\to H$ be the bijection that Duplicator selects at the next round. If $\operatorname{rk}(x_{1})=\operatorname{rk}(f^{\prime\prime}(x_{1}))=1$, then $f^{\prime\prime}(x_{1})\cdot f^{\prime}(x[2,\ldots,r+1])\neq f(x)$, since $\operatorname{rk}(f^{\prime\prime}(x_{1}))=1$ and $\operatorname{rk}(f^{\prime}(x[2,\ldots,q+1]))=q$, so their product has rank at most $q+1<\operatorname{rk}(f(x))$. In this case, Spoiler pebbles $x_{1}$ and wins immediately since the map on the pebbled elements $(x,x[2,\ldots,q+1],x_{1})\mapsto(f(x),f^{\prime}(x[2,\ldots,q+1]),f^{\prime\prime}(x_{1}))$ does not satisfy the winning condition for Spoiler in the Version I (and hence, the Version II) game.
@@
    If instead, $1=\operatorname{rk}(x_{1})<\operatorname{rk}(f^{\prime\prime}(x_{1}))$, Spoiler pebbles $x_{1}$ and wins with $k-1$ additional pebbles and $r$ additional rounds by assumption. Note that once $x_{1}\mapsto f^{\prime\prime}(x_{1})$ has been pebbled, Spoiler can reuse the pebble on $x$. So we only need $k-1$ additional pebbles rather than $k$ pebbles.
@@
  • Case 2: Suppose instead that the hypothesis of Case 1 is not satisfied. Then $\operatorname{rk}(y)\neq\operatorname{rk}(f^{\prime}(y))$ for some $y\in\langle C\rangle$ with $\operatorname{rk}(y)\leq q$. We will now do a binary-search-like procedure to eventually reach an element of $C$ that is mapped outside of $C$. To make this clearer, we introduce two additional parameter as in binary search. Let $lo=1$ and $hi=q+1$. In the next two rounds, Spoiler pebbles $y:=x[lo,\ldots,lo+\lceil(hi-lo)/2\rceil-1]$ and $z:=x[lo+\lceil(hi-lo)/2\rceil,\ldots,hi]$. Note that $\operatorname{rk}(y)=\lceil(hi-lo)/2\rceil$ and analogously for $\operatorname{rk}(z)$, with $\operatorname{rk}(x)=\operatorname{rk}(y)+\operatorname{rk}(z)$, for if $y$ or $z$ could be written as as shorter word in $C$, then so could $x$.
@@
    Let $f^{\prime\prime}:G\to H$ be the next bijection that Duplicator selects. If $f(x)\neq f^{\prime\prime}(y)f^{\prime\prime}(z)$ then Spoiler immediately wins. On the other hand, if $f(x)=f^{\prime\prime}(y)f^{\prime\prime}(z)$, then either
    $\operatorname{rk}(y)<\operatorname{rk}(f^{\prime\prime}(y))\qquad\text{ or }\qquad\operatorname{rk}(z)<\operatorname{rk}(f^{\prime\prime}(z)),$
    since $\operatorname{rk}(f(x))>\operatorname{rk}(x)=\operatorname{rk}(y)+\operatorname{rk}(z)$.
@@
    Without loss of generality, suppose that $\operatorname{rk}(y)<\operatorname{rk}(f^{\prime\prime}(y))$. Spoiler then updates $hi$ to $lo+\lceil(hi-lo)/2\rceil-1$, and as $y\mapsto f^{\prime\prime}(y)$ is already pebbled, Spoiler iterates on this strategy as in binary search, reusing the pebbles on $x$ and $z$ in the next iteration, with $y$ now playing the role that $x$ played in the first iteration. Thus, iterating this procedure uses at most $3$ pebbles (and we have $3\leq k$ by assumption). After $O(\log q)$ iterations, we reach the case where $hi=lo+1$, and we have pebbled an element $w\mapsto f^{\prime\prime\prime}(w)$ such that $1=\operatorname{rk}(w)<\operatorname{rk}(f^{\prime\prime\prime}(w))$, and there are two additional pebbles placed that can be reused.
@@
    By assumption, every element of $\langle C\rangle$ can be written as a word of length at most $d$ in the elements of $C$, so we have $q\leq d$, and thus only $O(\log d)$ rounds of the preceding iteration are required. (That $d\leq|\langle C\rangle|$ follows from Fact 2.2.)
@@
    Finally, since Spoiler has pebbled $w\in C$ mapping to an element outside of $C^{\prime}$, Spoiler implements the assumed strategy that identifies $C$ with $k$ pebbles and $r$ rounds. However, at the time Spoiler begins this strategy, two of the placed pebbles can be re-used, so the strategy only requires an additional $k-2$ pebbles. Plus the one pebble on $w$, for a total of $k-2+3=k+1$ pebbles.
Finally, we must handle the case of infinite rank. By symmetry, it suffices to show that Spoiler can win in the case when $\operatorname{rk}(x)=q+1$, but $\operatorname{rk}(f(x))=\infty$. In this case, the same argument starting from the third paragraph works mutatis mutandis, as $\operatorname{rk}(f(x))=\infty>q+1$. ∎
@@ 5 Direct products
Let $G=G_{1}\times\cdots\times G_{d}$ be a decomposition into indecomposable direct factors, let $k\geq 5$, and let $r:=r(n)$. If $G$ and $H$ are not distinguished by $(k,r+O(\log n))$-WL Version II, then there exist direct factors $H_{i}\leq H$ such that $H=H_{1}\times\cdots\times H_{d}$ such that for all $i\in[d]$, $G_{i}$ and $H_{i}$ are not distinguished by $(k-1,r)$-WL Version II.
@@ 5.1 Abelian and Semi-Abelian Case
    If there exist $x,y\in G$ such that $f([x,y])$ is not a commutator $[h,h^{\prime}]$ for any $h,h^{\prime}\in H$ (that is, $\operatorname{cw}(f([x,y]))>1$), then Spoiler can win with $3$ pebbles and $3$ rounds.
@@
    Spoiler pebbles $[x,y]\mapsto f([x,y])$. At the next two rounds, Spoiler pebbles $x,y$. Regardless of Duplicator’s choices, Spoiler wins.
@@
    Apply the Rank Lemma 4.3 to the set of commutators. By part (b), if $g\in[G,G]$ and $f(g)\not\in[H,H]$, then Spoiler can win with $3$ pebbles and $3$ rounds. So by the Rank Lemma 4.3, if $\operatorname{cw}(g)\neq\operatorname{cw}(f(g))$, then Spoiler can win with $4$ pebbles and $O(\log n)$ rounds.
@@
We adapt the proof of [BS22, Lemma 6.13] to control for rounds. Let $\widetilde{G}$ be a group such that $(4,O(1))$-WL Version II fails to distinguish $G$ and $\widetilde{G}$. By Lem. 5.7 and the subsequent discussion, we may assume that $Z(G)\cong Z(\widetilde{G})$ using $(2,2)$-WL Version II. As Abelian groups are direct products of their Sylow subgroups, it follows that $Z(G)$ and $Z(\widetilde{G})$ have isomorphic Sylow subgroups. Write $\widetilde{G}=\widetilde{H}\times\widetilde{A}$, where $\widetilde{A}$ is the maximal Abelian direct factor. As $Z(G)\cong Z(\widetilde{G})$, we write $Z$ for the Sylow $p$-subgroup of $Z(G)\cong Z(\widetilde{G})$. Consider the primary decomposition of $Z$:
@@
where $Z_{i}\cong(\mathbb{Z}/p^{i}\mathbb{Z})^{e_{i}}$, for $e_{i}\geq 1$. For each $i\in[m]$, there exist subgroups $H_{i}\leq Z(H)$ and $A_{i}\leq A$ such that $Z_{i}\cong H_{i}\times A_{i}$. Similarly, there exist $\widetilde{H_{i}}\leq\widetilde{H}$ and $\widetilde{A_{i}}\leq\widetilde{A}$ such that $Z_{i}\cong\widetilde{H_{i}}\times\widetilde{A_{i}}$. As $Z(G)\cong Z(\widetilde{G})$, we have that $H_{i}\times A_{i}\cong\widetilde{H_{i}}\times\widetilde{A_{i}}$. It suffices to show that each $A_{i}\cong\widetilde{A_{i}}$. As $H$ does not have any Abelian direct factors, we have by [BS22, Lemma 6.12] (reproduced as Lem. 5.11 above) that a central element $x$ of order $p^{i}$ splits from $G$ if and only if the projection of $x$ onto $A_{i}$, denoted $A_{i}(x)$, has order $p^{i}$. The same holds for $\widetilde{G}$ and the $\widetilde{A_{i}}$ factors. By Lem. 5.5, we may assume that Duplicator selects bijections $f:G\to\widetilde{G}$ such that if $x\in Z(G)$ splits from $Z_{i}$, then $f(x)$ splits from $f(Z_{i})$. Otherwise, Spoiler pebbles $x\mapsto f(x)$, and then following Lem. 5.5 (b), wins with $3$ additional pebbles (for a total of $4$ pebbles) and $3$ additional rounds (for a total of $4$ rounds). The result follows. ∎
@@
We follow the strategy of [BS22, Lemma 6.15]. First, we highlight our key changes. Their proofs has two parts: the fact that central $e$-th powers are detectable, and their Corollary 6.10. Using our Lem. 5.9 in place of their Corollary 6.10, we get 4 pebbles and $O(\log n)$ rounds, so all that is left to handle is central $e$-th powers. Suppose Duplicator selects a bijection $f:G\to H$ where $g=x^{e}$ for some $x\in Z(G)$ and $f(g)$ is not a central $e$th power. By Lem. 5.7 (a), Duplicator must map the center to the center or Spoiler can win with 2 pebbles in 2 rounds, so we need only handle the condition of being an $e$-th power. At the first round, Spoiler pebbles $g\mapsto f(g)$. At the next round, Spoiler pebbles $x$ and wins. Thus Spoiler can win with 2 pebbles in 2 rounds. ∎
@@
Let $G=H\times A$ and $\widetilde{G}=\widetilde{H}\times\widetilde{A}$, where $A$ and $\widetilde{A}$ are maximal Abelian direct factors. Let $k\geq 5$ and $r:=r(n)$. If $(k,r+O(\log n))$-WL Version II fails to distinguish $G$ and $\widetilde{G}$, then $(k-1,r)$-WL Version II fails to distinguish $H$ and $\widetilde{H}$.
@@
By Thm. 5.12, we may assume that $A\cong\widetilde{A}$ (otherwise, $(4,O(1))$-WL Version II distinguishes $G$ from $\widetilde{G}$). Consider the component-wise filtrations from the proof of [BS22, Lemma 6.15], $\{1\}=U_{0}\leq\cdots\leq U_{r}=Z(G)$ with respect to the decomposition $G=H\times A$ and $\{1\}=\widetilde{U_{0}}\leq\cdots\leq\widetilde{U_{r}}=Z(\widetilde{G})$ with respect to the decomposition $\widetilde{G}=\widetilde{H}\times\widetilde{A}$.
@@
Furthermore, Brachter & Schweitzer also showed in the proof of [BS22, Lemma 6.16] that Duplicator must map $H\times\{1\}$ to a system of representatives modulo $\{1\}\times\widetilde{A}$. Thus, Spoiler can restrict the game to $H\times\{1\}$. Now if $(k-1,r)$-WL Version II distinguishes $H$ and $\widetilde{H}$, then Spoiler can ultimately reach a configuration $((h_{1},1),\ldots,(h_{k-1},1))\mapsto(x_{1},a_{1}),\ldots,(x_{k-1},a_{k-1})$ such that the induced configuration over $(G/(\{1\}\times A),\widetilde{G}/(\{1\}\times\widetilde{A}))$ fulfills the winning condition for Spoiler. That is, considered as elements of $G/(\{1\}\times A)$ (resp., $\widetilde{G}/(\{1\}\times\widetilde{A})$), the map $(h_{1},\dotsc,h_{k-1})\mapsto(x_{1},\dotsc,x_{k-1})$ does not extend to an isomorphism. This implies that the pebbled map $((h_{1},1),\ldots,(h_{k-1},1))\mapsto((x_{1},a_{1}),\ldots,(x_{k-1},a_{k-1}))$ in the original groups (rather than their quotients) does not extend to an isomorphism. For suppose $f$ is any bijection extending the pebbled map. By the above, without loss of generality, $f$ maps $H\times\{1\}$ to a system of coset representatives of $\{1\}\times\widetilde{A}$, that is, if Duplicator can win, Duplicator can win with such a map. Let $\overline{f}$ be the induced bijection on the quotients $G/(\{1\}\times A)\to\tilde{G}/(\{1\}\times\tilde{A})$. Since the pebbled map on the quotients does not extend to an isomorphism, there is a word $w$ such that $\overline{f}(w(h_{1},\dotsc,h_{k-1}))\neq w(x_{1},\dotsc,x_{k-1})$. But then when we consider $f$ restricted to $H\times\{1\}$, we find that $f(w((h_{1},1),\dotsc,(h_{k-1},1)))=f((w(h_{1},\dotsc,h_{k-1}),1))\neq(w(x_{1},\dotsc,x_{k-1}),w(a_{1},\dotsc,a_{k-1}))$, because their $H$ coordinates are different. ∎

### v7

@@ 3.2 Coprime extensions with an $O(1)$-generated complement
Let $K$ be a group of order $n$ such that $K\not\cong G$. We will first show that if there is no action such that $K$ decomposes as $H\ltimes N$, then Spoiler can win with $O(1)$ pebbles and $O(1)$ rounds. Let $f:G\to K$ be the bijection that Duplicator selects. As $N\leq G$, as a subset, is uniquely determined by its orders—it is precisely the set of all elements in $G$ whose orders divide |N||N|—we may assume that $K$ has a normal Hall subgroup of size $|N|$. For first, if for some $n\in N$, $|n|\neq|f(n)|$, Spoiler can pebble $n\mapsto f(n)$ and win immediately. By reversing the roles of $K$ and $G$, it follows that $K$ must have precisely $|N|$ elements whose orders divide $|N|$. Second, suppose that those $|N|$ elements do not form a subgroup. Then there are two elements $x,y\in f(N)$ such that $xy\notin f(N)$. At the first round, Spoiler pebbles $a:=f^{-1}(x)\mapsto x$. Let $f^{\prime}:G\to K$ be the bijection Duplicator selects at the next round. As $K$ has precisely $|N|$ elements of order dividing $|N|$, we may assume that $f^{\prime}(N)=f(N)$ (setwise). Let $b\in N$ s.t. $f^{\prime}(b)=y$. Spoiler pebbles $b\mapsto y$. Now as $N$ is a group, $ab\in N$. However, as $f(a)f^{\prime}(b)\not\in f(N)$, $|ab|\neq|f(a)f^{\prime}(b)|$. So the map $(a,b)\mapsto(x,y)$ does not extend to an isomorphism. Spoiler now wins.
Now we have that $f(N)$ is a subgroup of $K$, and because it is the set of all elements of these orders, it is characteristic and thus normal. Suppose that $f(N)\not\cong N$. We have two cases: either $f(N)$ is not Abelian, or $f(N)$ is Abelian but $N\not\cong f(N)$. Suppose first that $f(N)$ is not Abelian. Let $x\in f(N)$ such that $x\not\in Z(f(N))$, and let $g:=f^{-1}(x)\in N$. Spoiler pebbles $g\mapsto x$. Let $f^{\prime}:G\to K$ be the bijection that Duplicator selects at the next round. We may again assume that $f^{\prime}(N)=f(N)$ (setwise), or Spoiler wins with two additional pebbles and two additional rounds. Now let $y\in f(N)$ such that $x,y]\neq 1$. Let $h\in G$ such that $f^{\prime}(h)=y$. Spoiler pebbles $h\mapsto y$. Now the map $(g,h)\mapsto(x,y)$ does not extend to an isomorphism, so Spoiler wins. Suppose instead that $f(N)$ is Abelian. As Abelian groups are determined by their orders, we have by the discussion in the first paragraph that Spoiler wins with $2$ pebbles and $2$ rounds.
@@ 4 A “rank” lemma
We say that $C\subseteq G$ is distinguished by the version $J$ pebble game with $k$ pebbles and $r$ rounds if for all $g\in C$ and $g^{\prime}\notin C$, Spoiler can win this pebble game, played on $G$ and $G$, starting from the configuration $g\mapsto g^{\prime}$. In terms of WL colorings, this is equivalent to: after $r$ rounds of $(k+1)$-WL Version $J$ coloring, the colors of all tuples containing elements of $C$ are distinct from the colors of all tuples not containing elements of $C$. In particular, if $C\subseteq G$ is distinguished by $(k+1,r)$-WL, then for any group $H$ of the same order as $G$, we define $C_{H}$ to be the set of those elements in $H$ whose $(k+1,r)$-WL colors are among the $(k+1,r)$-WL colors of $C$. It follows that in the Version $J$, $k$-pebble game on $G$ and $H$, Duplicator must choose bijections $f\colon G\to H$ that map $C$ to $C_{H}$, or lose within an additional $r$ rounds.
@@
Let $k\geq 3$, and $J\in\{I,II\}$. Suppose $C\subseteq G$ is distinguished (as defined above) by the Version $J$ pebble game with $k$ pebbles and $r$ rounds; let $C_{H}$ be the corresponding subset of $H$ as above. In the pebble game played on $G$ and another group $H$, suppose that at the end of a round there is a pebble on $g\mapsto h$ such that $\operatorname{rk}_{C}(g)\neq\operatorname{rk}_{C_{H}}(h)$. Then Spoiler can win with $k$ additional pebbles (beyond the one on $g\mapsto h$) and $r+\log d+O(1)$ additional rounds, where $d=\operatorname{rk}_{C}(g)\leq\text{diam}(\text{Cay}(\langle C\rangle,C))\leq|\langle C\rangle|\leq|G|$.
Our primary uses of this lemma in this paper are to show that if $C$ is distinguished by $(k,r)$-WL, then $\langle C\rangle$ is distinguished by $(k,r+\log n+O(1))$-WL. However, the preservation of $C$-rank itself, rather than merely the subgroup generated by $C$, seems potentially useful for future applications. In particular, Lem. 4.3 shows that WL can identify verbal subgroups in $O(\log n)$ rounds, provided WL can readily identify each word.
One sees that the version of the Rank Lemma stated in the Introduction (Lemma 1.13) follows immediately from the above: if Duplicator chooses a bijection $f$ that does not preserve the rank of $g$, Spoiler’s first move is to pebble $g\mapsto f(g)$, and then Lemma 4.3 applies (note the $k+1$ in the statement of Lemma 1.13, compared to $k$ in Lemma 4.3, is to account for this first pebble).
@@
Throughout, we will use $\operatorname{rk}(x)$ to denote $\operatorname{rk}_{C}(x)$ if $x\in G$, and $\operatorname{rk}_{C_{H}}(x)$ if $x\in H$.
Without loss of generality (by swapping the roles of $G$ and $H$ if needed), we may assume that $\operatorname{rk}(g)<\operatorname{rk}(h)$. If $\operatorname{rk}(g)=1$, then by assumption, since $C=\{g\in G:\operatorname{rk}(g)=1\}$, Spoiler wins with $k$ pebbles in $r$ rounds.
@@
On the other hand, if $\operatorname{rk}(g)>1$, let $f:G\to H$ be the bijection that Duplicator selects at the next round. For brevity let $q=\operatorname{rk}(g)$. Note that this case works even if $\operatorname{rk}(h)=\infty$, for then by our assumption that $\operatorname{rk}(g)<\operatorname{rk}(h)$, we still have $q=\operatorname{rk}(g)$ is finite. Now write $g=g_{1}\cdots g_{q}$, where for each $i$, $g_{i}\in C$. For $1\leq i\leq j\leq q$, write $g[i,\ldots,j]:=g_{i}\cdots g_{j}$. We consider the following cases.
@@
  • Case 1: Suppose first that for all $x\in G$ with $\operatorname{rk}(x)<q$, we have $\operatorname{rk}(x)=\operatorname{rk}(f(x))$. In this case, Spoiler pebbles $g[2,\ldots,q]\mapsto f(g[2,\ldots,q])$. Let $f^{\prime}:G\to H$ be the bijection that Duplicator selects at the next round; Spoiler pebbles $g_{1}\mapsto f^{\prime}(g_{1})$.
    If $\operatorname{rk}(g_{1})=\operatorname{rk}(f^{\prime}(g_{1}))=1$, then $f^{\prime}(g_{1})\cdot f(g[2,\ldots,r+1])\neq h$, since $\operatorname{rk}(f^{\prime}(g_{1}))=1$ and $\operatorname{rk}(f(g[2,\ldots,q]))=q-1$, so their product has rank at most $q=\operatorname{rk}(g)<\operatorname{rk}(h)$. In this case, Spoiler wins immediately since the map on the pebbled elements $(g,g[2,\ldots,q],g_{1})\mapsto(h,f(g[2,\ldots,q]),f^{\prime}(g_{1}))$ does not satisfy the winning condition for Spoiler in the Version I (and hence, the Version II) game.
@@
    If instead, $1=\operatorname{rk}(g_{1})<\operatorname{rk}(f^{\prime}(g_{1}))$, then Spoiler wins with k1k-1 additional pebbles and $r$ additional rounds by assumption. Note that once $g_{1}\mapsto f^{\prime}(g_{1})$ has been pebbled, Spoiler can reuse the pebble on $g$; this is why we only need k1k-1 additional pebbles in this part rather than $k$ additional pebbles.
@@
  • Case 2: Suppose instead that the hypothesis of Case 1 is not satisfied. Then there exists some $x\in\langle C\rangle$ with $\operatorname{rk}(x)\neq\operatorname{rk}(f(x))$ and $\operatorname{rk}(x)<q=\operatorname{rk}(g)$. We will now do a binary-search-like procedure to eventually reach an element of $C$ that is mapped outside of $C$. To make this clearer, we introduce two additional parameters as in binary search. Let $lo=1$ and $hi=q$. In the next two rounds, Spoiler pebbles $x:=g[lo,\ldots,lo+\lceil(hi-lo)/2\rceil-1]$ and $y:=g[lo+\lceil(hi-lo)/2\rceil,\ldots,hi]$. Note that $\operatorname{rk}(x)=\lceil(hi-lo)/2\rceil$ and $\operatorname{rk}(y)=(hi-lo)-\lceil(hi-lo)/2\rceil+1$, with $\operatorname{rk}(g)=\operatorname{rk}(x)+\operatorname{rk}(y)$, for if $x$ or $y$ could be written as a shorter word in $C$, then so could $g$.
@@
    Let $f^{\prime}:G\to H$ be the next bijection that Duplicator selects. If $h\neq f^{\prime}(x)f^{\prime}(y)$ then Spoiler immediately wins. On the other hand, if $h=f^{\prime}(x)f^{\prime}(y)$, then either
    $\operatorname{rk}(x)<\operatorname{rk}(f^{\prime}(x))\qquad\text{ or }\qquad\operatorname{rk}(y)<\operatorname{rk}(f^{\prime}(y)),$
    since $\operatorname{rk}(h)>\operatorname{rk}(g)=\operatorname{rk}(x)+\operatorname{rk}(y)$.
@@
    Without loss of generality, suppose that $\operatorname{rk}(x)<\operatorname{rk}(f^{\prime}(x))$. Spoiler then updates $hi$ to $lo+\lceil(hi-lo)/2\rceil-1$, and as $x\mapsto f^{\prime}(x)$ is already pebbled, Spoiler iterates on this strategy as in binary search, reusing the pebbles on $g$ and $y$ in the next iteration, with $x$ now playing the role that $g$ played in the first iteration. Thus, iterating this procedure uses at most $3$ pebbles (and we have $3\leq k$ by assumption). After $\log q+O(1)$ iterations, we reach the case where $hi=lo+1$, and we have pebbled an element $z\mapsto f^{\prime\prime}(z)$ such that $1=\operatorname{rk}(z)<\operatorname{rk}(f^{\prime\prime}(z))$, and there are two additional pebbles placed that can be reused.
@@
    By assumption, every element of $\langle C\rangle$ can be written as a word of length at most $d$ in the elements of $C$, so we have $q\leq d$, and thus only $\log d+O(1)$ rounds of the preceding iteration are required. (That $d\leq|\langle C\rangle|$ follows from Fact 2.2.)
@@
    Finally, since Spoiler has pebbled $z\in C$ mapping to an element outside of $C_{H}$, Spoiler implements the assumed strategy that identifies $C$ with $k$ pebbles and $r$ rounds. However, at the time Spoiler begins this strategy, two of the placed pebbles can be re-used, so the strategy only requires an additional k2k-2 pebbles, for a total of $k$ additional pebbles beyond the one that we started with on $g\mapsto h$. (We do not need to count +1+1 for the pebble on $z$, since the definition of “distinguished” includes “starting from the position $z\mapsto f^{\prime\prime}(z)$].”) ∎
@@ 5 Direct products
Let $G=G_{1}\times\cdots\times G_{d}$ be a decomposition into indecomposable direct factors, let $k\geq 5$, and let $r:=r(n)$. If $G$ and $H$ are not distinguished by $(k,r+O(\log n))$-WL Version II, then there exist direct factors $H_{i}\leq H$ such that $H=H_{1}\times\cdots\times H_{d}$ such that for all $i\in[d]$, $G_{i}$ and $H_{i}$ are not distinguished by (k1,r)(k-1,r)-WL Version II.
@@ 5.1 Abelian and Semi-Abelian Case
    If there exist $x,y\in G$ such that $f([x,y])$ is not a commutator $h,h^{\prime}]$ for any $h,h^{\prime}\in H$ (that is, $\operatorname{cw}(f([x,y]))>1$), then Spoiler can win with $3$ pebbles and $3$ rounds.
@@
    Spoiler pebbles $x,y]\mapsto f([x,y])$. At the next two rounds, Spoiler pebbles $x,y$. Regardless of Duplicator’s choices, Spoiler wins.
@@
    Apply the Rank Lemma 4.3 to the set of commutators. Suppose that there exist $x,y\in G$ such that $g=[x,y]$, but for all $x^{\prime},y^{\prime}\in H$, $f(g)\neq[x^{\prime},y^{\prime}]$. Then by part (b), Spoiler can win with $3$ pebbles and $3$ rounds. So by the Rank Lemma 1.13, if $\operatorname{cw}(g)\neq\operatorname{cw}(f(g))$, then Spoiler can win with $4$ pebbles and $O(\log n)$ rounds.
@@
[…]
@@
[…]
@@
We follow the strategy of [BS22, Lemma 6.15]. First, we highlight our key changes. Their proof has two parts: the fact that central $e$-th powers are detectable, and their Corollary 6.10. Using our Lem. 5.9 in place of their Corollary 6.10, we get 4 pebbles and $O(\log n)$ rounds, so all that is left to handle is central $e$-th powers. Suppose Duplicator selects a bijection $f:G\to H$ where $g=x^{e}$ for some $x\in Z(G)$ and $f(g)$ is not a central $e$th power. By Lem. 5.7 (a), Duplicator must map the center to the center or Spoiler can win with 2 pebbles in 2 rounds, so we need only handle the condition of being an $e$-th power. At the first round, Spoiler pebbles $g\mapsto f(g)$. At the next round, Spoiler pebbles $x$ and wins. Thus Spoiler can win with 2 pebbles in 2 rounds. ∎
@@
Let $G=H\times A$ and $\widetilde{G}=\widetilde{H}\times\widetilde{A}$, where $A$ and $\widetilde{A}$ are maximal Abelian direct factors. Let $k\geq 5$ and $r:=r(n)$. If $(k,r+O(\log n))$-WL Version II fails to distinguish $G$ and $\widetilde{G}$, then (k1,r)(k-1,r)-WL Version II fails to distinguish $H$ and $\widetilde{H}$.
@@
[…]
@@
[…]